OpenStudy (bookie2cute):

(p + q = 1, p2 + 2pq + q2 = 1) What the answer is ??/

OpenStudy (mathstudent55):

p + q = 1 p2 + 2pq + q2 = 1 Factor the second equation: (p + q)(p + q) = 1 From the first equation, we know that p + q = 1, so substitute 1 in for p + q once: 1(p + q) = 1 p + q = 1 q = 1 - p This is true for every p and q such that q = 1 - p